can you help me with this problem?
Two consecutive numbers can be represented by x and x+1. The product of two consecutive numbers is equal to 35 more than 11 times the sum of the numbers. Write a quadratic equation represent the situation.
I got (x)(x+1)= and that's it. Can you help me by showing your work, not so I can copy, but because I have absolutely no idea how to solve this problem. Thank you :)

Answer by Alan3354(69443)

Two consecutive numbers can be represented by x and x+1. The product of two consecutive numbers is equal to 35 more than 11 times the sum of the numbers. Write a quadratic equation represent the situation.
I got (x)(x+1)= and that's it.
------------------
The sum of the numbers is x + x+1
11 times that is 11*(x + x+1)
Then add 35
x(x+1) = 11*(x + x+1) + 35